OpenStudy (lilah):

Write the formulas that may be used to solve the problems below, then solve. A. How long will it take Karen to drive 975 miles if her average speed is 65 miles per hour? B. A rectangular pool has a tile border around its perimeter that measures 66 feet. If the length of the pool is 18 feet, what is the width of the pool? C. A current of 0.05 amps is driven by a voltage of 0.6 volts. What is the resistance measured in ohms? D. A thermometer shows a patient's temperature to be 39 degrees celsius. What is the patient's temperature in degrees fahrenheit?

OpenStudy (lilah):

@pooja195

OpenStudy (lilah):

@whpalmer4

OpenStudy (lilah):

@sweetburger

MARC (marc_d):

question A\[Average~speed=\frac{ distance }{ time~taken }\]given that the distance travelled=975 miles..and the average speed=65 miles/per hour..plug in all the value to find the time taken travelled by Karen..

MARC (marc_d):

\[time~taken=\frac{ distance }{ average~speed }\]

MARC (marc_d):

\[time~taken=\frac{ 975miles }{ 65miles~h^{-1} }\]

MARC (marc_d):

\[time~taken=\frac{ 15\cancel{975miles} }{ 1\cancel{65miles}h^{-1} }\]

MARC (marc_d):

\[time~taken=?\]

MARC (marc_d):

question B\[perimeter~of~a~rectangular~pool=2(l+w)\]where l=length w=width given that the perimeter of a rectangular pool=66 feet..and the length of the pool=18 feet..plug in all the value to find the width of the rectangular pool..

MARC (marc_d):

\[perimeter=66\]

MARC (marc_d):

\[2(l+w)=66\]

MARC (marc_d):

\[2[(18)+w]=66\]

MARC (marc_d):

divide both sides by 2

MARC (marc_d):

\[18+w=33\]

MARC (marc_d):

subtract both sides by 18

MARC (marc_d):

\[w=33-18\]

MARC (marc_d):

\[w=?\]

MARC (marc_d):

question C\[R=\frac{ V }{ I }\]where R=resistance V=voltage I=current given that current=0.05A..and voltage=0.6V...plug inall the values to find the value of R..

MARC (marc_d):

\[R=\frac{ 0.6 }{ 0.05 }\]

MARC (marc_d):

\[R=?\]

MARC (marc_d):

question D\[T_{(F)}=T_{(C)}\times1.8+32\]where T_(F)=Fahrenheit T_(C)=Celsius given that the patient's temperature=39 dgree celsius=T_(C)..plug in all the values to find the temperature in degrees fahrenheit..

MARC (marc_d):

\[T_{(F)}=39\times1.8+32\]

MARC (marc_d):

\[T_{(F)}=70.2+32\]

MARC (marc_d):

\[T_{(F)}=?\]

MARC (marc_d):

solve it! @Lilah